Solve the following equation by clearing fractions. Please show all of your work.
2/7+x=x/3-3

Answer by Alan3354(69443)

2/7+x=x/3-3
Multiply by 7
2 + 7x = (7x/3) - 21
Multiply by 3
6 + 21x = 7x - 63
6 + 14x = -63
14x = -69
x = -69/14